OpenStudy (anonymous):

I need help expressing y as a function of x. 2ln y=-1/2ln x+1/3ln(x^2+1)+ln c

zepdrix (zepdrix):

Hey Squirrel :) Welcome to OpenStudy!

zepdrix (zepdrix):

y a function of x? Hmmm... let's see...

zepdrix (zepdrix):

\[\Large\rm 2\ln y=-\frac{1}{2}\ln x+\frac{1}{3}\ln(x^2+1)+\ln c\]Let's divide both sides by 2 to get isolate our `ln y`,\[\Large\rm \ln y=-\frac{1}{4}\ln x+\frac{1}{6}\ln(x^2+1)+\frac{1}{2}\ln c\]

zepdrix (zepdrix):

We would like to condense this down to one or two logs on the right side, then we can do some exponentiation.

zepdrix (zepdrix):

So we need to apply some log rules.

zepdrix (zepdrix):

\[\Large\rm \ln y=\ln\left(\frac{1}{\sqrt[4]{x}}\right)+\frac{1}{6}\ln(x^2+1)+\frac{1}{2}\ln c\]Do you understand what I did in that first step? Or should I slow it down a lil bit?

OpenStudy (anonymous):

what log rules did you apply?

zepdrix (zepdrix):

\[\Large\rm a \log(b)=\log(b^a)\]Since the thing out front was a fraction, I rewrote it as a root.\[\Large\rm -\frac{1}{4}\ln(x)=\ln\left(x^{-1/4}\right)=\ln\left(\frac{1}{x^{1/4}}\right)=\ln\left(\frac{1}{\sqrt[4]{x}}\right)\]

OpenStudy (anonymous):

oh ok thank you now that step is clear

OpenStudy (anonymous):

do we ever cancel the ln?

zepdrix (zepdrix):

There is a way we can deal with the logs, but not yet. We first to get it into this form: ln(y) = ln(stuff) We would like a single log on the right side.

zepdrix (zepdrix):

After you apply that same log rule to the other two terms, you get something like this:\[\Large\rm \ln y=\ln\left(\frac{1}{\sqrt[4]{x}}\right)+\ln\left(\sqrt[6]{x^2+1}\right)+\ln\sqrt{c}\]

zepdrix (zepdrix):

Then we need to apply this rule:\[\Large\rm \log(a)+\log(b)=\log(ab)\]

OpenStudy (anonymous):

is that the product one?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

and we combine it?

zepdrix (zepdrix):

Yes, so combining the first two terms should give us something like this,\[\Large\rm \ln y=\ln\left(\frac{\sqrt[6]{x^2+1}}{\sqrt[4]{x}}\right)+ln\sqrt{c}\]Yah?

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

we cannot reduce the 6 and 4?

zepdrix (zepdrix):

No. The numerator and denominator do not have the same base.

OpenStudy (anonymous):

ok is there anything i can do here in terms of reducing?

zepdrix (zepdrix):

So apply that log rule one more time, \[\Large\rm \ln y=\ln\left(\frac{\sqrt[6]{x^2+1}~\sqrt c}{\sqrt[4]{x}}\right)\]

zepdrix (zepdrix):

And now it's in the form that we wanted:\[\Large\rm \ln y=\ln(stuff)\]

zepdrix (zepdrix):

We can exponentiate each side, Or maybe easier yet, just recognize that each side is inside of a log, and each log is equal to one another, then the insides of those logs have to be equal as well.\[\Large\rm y=\frac{\sqrt[6]{x^2+1}~\sqrt{c}}{\sqrt[4]{x}}\]

OpenStudy (anonymous):

so that gets rid of the ln?

zepdrix (zepdrix):

Yes. If you prefer to be a little bit more careful about it, you can exponentiate each side:\[\Large\rm \ln y=\ln(stuff)\]\[\Large\rm e^{\ln y}=e^{\ln(stuff)}\]Since the exponential and log are inverse operations of one another they essentially "undo" one another,\[\Large\rm y=stuff\]

OpenStudy (anonymous):

ok makes sense, btw thank you thank you this has been so helpful

OpenStudy (anonymous):

is that the final answer?